### v1

# An accurate mean-field equation for voter model dynamics on scale-free networks

Abstract
Understanding the emergent macroscopic behavior of dynamical systems on networks is a crucial but challenging task. One of the simplest and most effective methods to construct a reduced macroscopic model is given by mean-field theory. The resulting approximations perform well on dense and homogeneous networks but poorly on scale-free networks, which, however, are more realistic in many applications. In this paper, we introduce a modified version of the mean-field approximation for voter model dynamics on scale-free networks. This approach achieves a significantly smaller approximation error than standard methods without increasing dimensionality.
I Introduction
The voter model [1, 2] and its many extensions (see, e.g., [3, 4, 5] for overviews) are among the most popular models for studying spreading or infection processes on networks. In this paper, we focus on a general variant, the continous-time noisy voter model (CNVM) [6, 7, 8]. Given a simple graph $G$ on $N$ nodes, each node $i\in\{1,\dots,N\}$ has a state $x_{i}\in\{0,\dots,M-1\}=:[M]$. As the voter model is often applied to opinion dynamics, we refer to $x_{i}$ as the opinion of node $i$. Each opinion $x_{i}$ evolves over time according to a continuous-time Markov chain: if $x_{i}$ is in state $a\in[M]$, it transitions to a different state $b\in[M]$ at the rate
$r_{a,b}\frac{d_{i,b}}{d_{i}}+\tilde{r}_{a,b},$ (1)
where $d_{i,b}$ denotes the number of neighbors of node $i$ with opinion $b$, $d_{i}$ is the degree of node $i$, and $r_{a,b},\tilde{r}_{a,b}\geq 0$ are model parameters. The first term in (1) describes imitation: the more neighbors of node $i$ have opinion $b$, the higher the transition rate to $b$. The second term $\tilde{r}_{a,b}$ is independent of the neighborhood and induces explorative behavior, which can also be interpreted as noise.
An important observable for such discrete-state systems is the vector of shares $c\in[0,1]^{M}$, $\lVert c\rVert_{1}=1$, where $c_{a}$ describes the fraction of nodes $i$ that have opinion $x_{i}=a$. Using mean-field theory [9], it has been shown that, in the large-population limit $N\to\infty$, the trajectory $c(t)$ of time-dependent shares converges to the solution of a mean-field equation (MFE), provided that the underlying network is sufficiently homogeneous and dense [6, 10, 11, 12]. For scale-free networks—e.g., generated by the Barabási–Albert (BA) model [13] with preferential attachment parameter mm—however, this classical mean-field approximation performs poorly [14, 15, 16]. The main contribution of this paper is the derivation and validation of a modified mean-field equation (mMFE) that accurately captures the macroscopic behavior of the CNVM on BA networks.
It is well known [6] that for the CNVM the MFE is a simple system of ordinary differential equations
$\frac{\text{d}}{\text{d}t}c(t)=\sum_{\begin{subarray}{c}a,b\in[M]\\ a\neq b\end{subarray}}c_{a}(t)\big{(}r_{a,b}\,c_{b}(t)+\tilde{r}_{a,b}\big{)}(e_{b}-e_{a}),$ (2)
where $e_{a}$ denotes the $a$-th unit vector in $\mathbb{R}^{M}$. To achieve an accurate approximation for BA networks, we introduce two modifications. First, the imitation rates $r_{a,b}$ are multiplied by a constant correlation factor $0<\alpha\leq 1$; second, the shares $c\in[0,1]^{M}$ are replaced with the degree-weighted shares $\bar{c}\in[0,1]^{M}$, defined as the average over the states weighted by node degrees. The dynamics of these weighted shares is then approximated by the mMFE
$\frac{\text{d}}{\text{d}t}\bar{c}(t)=\sum_{a\neq b}\bar{c}_{a}(t)\big{(}\alpha\,r_{a,b}\,\bar{c}_{b}(t)+\tilde{r}_{a,b}\big{)}(e_{b}-e_{a}).$ (3)
The significance of the degree-weighted shares is already known in some contexts [17, 18], and they were specifically determined to be excellent collective variables for the CNVM on BA networks using a data-driven method developed by the authors [19].
In this paper, we derive the above mMFE (3), motivating why these modifications are necessary for BA networks, and show how to obtain the correlation factor $\alpha$. We analyze the quality and robustness of the mMFE depending on the model parameters $r$ and $\tilde{r}$, using a theoretical error analysis as well as numerical studies. We show that the mMFE approximates the degree-weighted shares exceptionally well when the imitation rates $r_{a,b}$ are similarly large and there is moderate noise $\tilde{r}$. Moreover, in this setting the mMFE is robust in the sense that $\alpha$ does not depend on the specific model parameters and initial state. In other parameter regimes, the approximation quality degrades slightly. We show that the mMFE outperforms standard methods like MFE and pair-approximation (PA) [4, 15] significantly. Another interesting result is that the correlation factor $\alpha$ increases towards $1$ with increasing preferential attachment parameter $m$ of the BA network.
Accompanying code for reproducing the numerical experiments presented in this paper is provided at [20].
II The modified mean-field equation
In this section, we derive the modified mean-field equation (mMFE) from the evolution equation of the first-order moments. We have to rely on two assumptions, the correlation assumption and the interchangeability assumption, which are explained in more detail later. For simplicity, we consider the case of $M=2$ opinions. The general case is analogous, and is treated in the appendix A. In the binary-state case, it is sufficient to only work with the (expected) weighted share $\bar{c}_{1}$ of opinion $1$, which we will denote as $\bar{c}=\bar{c}_{1}$ for simplicity. It can be defined as
$\bar{c}:=\mathbb{E}\Big{[}\frac{1}{\sum_{i}d_{i}}\sum_{i\in[N]}d_{i}x_{i}\Big{]}=\frac{1}{\sum_{i}d_{i}}\sum_{i\in[N]}d_{i}\mathbb{E}[x_{i}],$ (4)
where expectation is taken with respect to stochasticity in the dynamical evolution and possibly uncertainty in the initial state. Recall that in the CNVM the rate at which node $i$ in state $x_{i}=0$ transitions to state $1$ is given by
$r_{0,1}\frac{1}{d_{i}}\sum_{j\in[N]}A_{i,j}x_{j}+\tilde{r}_{0,1},$ (5)
where $A$ denotes the adjacency matrix of the network. Hence, the evolution of the first-order moment $\mathbb{E}[x_{i}]$ is
$\frac{\text{d}}{\text{d}t}\mathbb{E}[x_{i}]=$
$\mathbb{P}(x_{i}=0)\ \mathbb{E}\Big{[}\frac{r_{0,1}}{d_{i}}\sum_{j}A_{i,j}x_{j}+\tilde{r}_{0,1}\mid x_{i}=0\Big{]}\$-
$\mathbb{P}(x_{i}=1)\ \mathbb{E}\Big{[}\frac{r_{1,0}}{d_{i}}\sum_{j}A_{i,j}(1-x_{j})+\tilde{r}_{1,0}\mid x_{i}=1\Big{]}.$ (6)
By inserting the equality $\mathbb{P}(x_{i}=0)\,\mathbb{E}[x_{j}\mid x_{i}=0]=\mathbb{E}[(1-x_{i})x_{j}]$, this implies
$\frac{\text{d}}{\text{d}t}\mathbb{E}[x_{i}]$ $=\frac{r_{0,1}}{d_{i}}\sum_{j}A_{i,j}\mathbb{E}[(1-x_{i})x_{j}]+\tilde{r}_{0,1}\mathbb{E}[1-x_{i}]$
$-\frac{r_{1,0}}{d_{i}}\sum_{j}A_{i,j}\mathbb{E}[x_{i}(1-x_{j})]-\tilde{r}_{1,0}\mathbb{E}[x_{i}].$ (7)
Hence, it follows that
$\frac{\text{d}}{\text{d}t}\bar{c}=\frac{1}{\sum_{i}d_{i}}\sum_{i}d_{i}\frac{\text{d}}{\text{d}t}\mathbb{E}[x_{i}]=$ (8)
$\frac{1}{\sum_{i}d_{i}}\Big{(}r_{0,1}\sum_{i,j}A_{i,j}\mathbb{E}[(1-x_{i})x_{j}]+\tilde{r}_{0,1}\sum_{i}d_{i}\mathbb{E}[1-x_{i}]$
$\quad-r_{1,0}\sum_{i,j}A_{i,j}\mathbb{E}[x_{i}(1-x_{j})]-\tilde{r}_{1,0}\sum_{i}d_{i}\mathbb{E}[x_{i}]\Big{)}.$ (9)
We break up the second-order moments via the (weak) correlation assumption
$\sum_{i,j}A_{i,j}\mathbb{E}[(1-x_{i})x_{j}]\overset{!}{=}\sum_{i,j}A_{i,j}\,\alpha\,\mathbb{E}[1-x_{i}]\mathbb{E}[x_{j}]$ (10)
for some constant correlation factor $0<\alpha<1$. This implies that, on average, it is more likely to find two neighbors in the same state than if they were stochastically independent. Given this assumption, we have
$\frac{\text{d}}{\text{d}t}\bar{c}=\frac{1}{\sum_{i}d_{i}}\Big{(}$
$\quad r_{0,1}\sum_{i,j}A_{i,j}\,\alpha\,\mathbb{E}[1-x_{i}]\mathbb{E}[x_{j}]+\tilde{r}_{0,1}\sum_{i}d_{i}\mathbb{E}[1-x_{i}]$
$-r_{1,0}\sum_{i,j}A_{i,j}\,\alpha\,\mathbb{E}[x_{i}]\mathbb{E}[1-x_{j}]-\tilde{r}_{1,0}\sum_{i}d_{i}\mathbb{E}[x_{i}]\Big{)}.$ (11)
Next, we apply the (weak) interchangeability assumption
$\sum_{i,j}A_{i,j}\,\mathbb{E}[1-x_{i}]\mathbb{E}[x_{j}]\overset{!}{=}\sum_{i,j}A_{i,j}\,(1-\bar{c})\bar{c},$ (12)
which is a weaker form of stating that all nodes have the same average state $\bar{c}$, and are therefore interchangeable. Inserting yields the mMFE for $M=2$ opinions
$\frac{\text{d}}{\text{d}t}\bar{c}\ =$ $\quad\alpha\,r_{0,1}(1-\bar{c})\bar{c}+\tilde{r}_{0,1}(1-\bar{c})$
$-\alpha\,r_{1,0}\,\bar{c}(1-\bar{c})-\tilde{r}_{1,0}\,\bar{c}.$ (13)
An analogous derivation yields the mMFE for an arbitrary number $M$ of opinions, as defined in (3), see appendix A for details.
It is clear that, due to possible violation of the two assumptions, the solution of the mMFE (3) does not exactly match the actual weighted shares $\bar{c}$. To distinguish the two, we will continue to refer to the true weighted shares as $\bar{c}$ and denote the solution of the mMFE as $\bar{c}_{\text{{\tiny mMFE}}}$.
II.1 Relation to classical mean-field theory
The two assumptions (10) and (12) are closely related to those used in classical mean-field theory, but adapted to reflect the properties of scale-free networks. In the classical theory, it is assumed that all node degrees grow unbounded as the number of nodes increases and hence any two neighbors become stochastically independent, i.e., choosing $\alpha=1$ in (10). In scale-free networks, however, a large portion of nodes have finite low degree, regardless of how large the network grows. As a result, the states of neighboring nodes do not become stochastically independent, but instead remain positively correlated. The important observation here is that the average correlation between neighbors can, with some limitations outlined later, be well captured by a constant factor $\alpha<1$.
Moreover, in the classical theory it is typically argued that nodes become interchangeable in the mean-field limit, so that the expected state of every node is simply given by the average state, e.g., $\mathbb{E}[x_{i}]=c$ for all $i$ in the binary-state case ($M=2$). We observe numerically that, for scale-free networks, this interchangeability is also valid on average, although the expected state of each node is related to the degree-weighted average state $\bar{c}$ instead, as described in assumption (12). As mentioned earlier, this finding was inspired by the authors’ work [19], which revealed computationally that the weighted shares are excellent collective variables for the CNVM on BA networks.
Further verification of the above claims will be given in the following sections.
II.2 Error analysis
Again, for simplicity, we assume the case of $M=2$ opinions. We define the error terms $\delta_{\text{cor}},\delta_{\text{int}}\in[-1,1]$ of the two assumptions (10) and (12) as
$\delta_{\text{cor}}$ $:=\frac{1}{\sum_{i}d_{i}}\sum_{i,j}A_{i,j}\Big{(}$ (14)
$\qquad\qquad\alpha\,\mathbb{E}[(1-x_{i})]\mathbb{E}[x_{j}]-\mathbb{E}[(1-x_{i})x_{j}]\Big{)}$
$\delta_{\text{int}}$ $:=\frac{1}{\sum_{i}d_{i}}\sum_{i,j}A_{i,j}\Big{(}$ (15)
$\qquad\qquad(1-\bar{c})\bar{c}-\mathbb{E}[(1-x_{i})]\mathbb{E}[x_{j}]\Big{)}.$
Note that $\delta_{\text{cor}}(t;\alpha)$ depends on time and on the factor $\alpha$, whereas $\delta_{\text{int}}(t)$ only depends on time. Let $F$ denote the right-hand side of the mMFE (II), i.e.,
$\frac{\text{d}}{\text{d}t}\bar{c}_{\text{{\tiny mMFE}}}=F(\bar{c}_{\text{{\tiny mMFE}}}).$ (16)
Recall that the evolution of the exact weighted shares $\bar{c}$ is given by the first-order moment equation (8). Inserting the error terms $\delta_{\text{cor}}$ and $\delta_{\text{int}}$ into (8) yields
$\frac{\text{d}}{\text{d}t}\bar{c}=F(\bar{c})+(r_{1,0}-r_{0,1})(\delta_{\text{cor}}+\alpha\,\delta_{\text{int}}).$ (17)
Moreover, note that $F$ is Lipschitz with constant $K(\alpha):=\alpha\lvert r_{1,0}-r_{0,1}\rvert+\lvert\tilde{r}_{1,0}-\tilde{r}_{0,1}\rvert$. Thus, by Gronwall’s inequality we obtain an upper bound for the error
$\lvert\bar{c}(t)-\bar{c}_{\text{{\tiny mMFE}}}(t;\alpha)\rvert\leq$ (18)
$\quad e^{K(\alpha)t}\int_{0}^{t}\big{\lvert}(r_{1,0}-r_{0,1})(\delta_{\text{cor}}(s;\alpha)+\alpha\,\delta_{\text{int}}(s))\big{\rvert}\text{d}s.$
Since $\lvert\delta_{\text{cor}}+\alpha\,\delta_{\text{int}}\rvert\leq 2$, this automatically implies a small error if $r_{1,0}\approx r_{0,1}$ and $\tilde{r}_{1,0}\approx\tilde{r}_{0,1}$. However, if these rate differences are large, the overall error can also become large, depending on $\delta_{\text{cor}}$ and $\delta_{\text{int}}$—that is, on how well the two assumptions (10) and (12) of the derivation hold.
Let the number of opinions $M$ be arbitrary again. To assess the quality of the mMFE, we employ the normalized $L_{1}$-error over an interval $[0,T]$
$L_{1}(T;\alpha):=\frac{1}{T}\int_{0}^{T}\frac{1}{2}\lVert\bar{c}(t)-\bar{c}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\text{d}t.$ (19)
(Note that $\frac{1}{2}\lVert\bar{c}(t)-\bar{c}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\in[0,1]$ is the total variation distance between discrete probability distributions [21].) Then, choosing the factor
$\alpha^{*}:=\operatorname*{argmin}_{\alpha}L_{1}(T;\alpha)$ (20)
yields the optimal fit of $\bar{c}_{\text{{\tiny mMFE}}}$ to $\bar{c}$ on the interval $[0,T]$.
II.3 Parameter regimes
Figure 1: Errors $\delta_{\text{cor}}(t;\alpha^{*})$, $\delta_{\text{int}}(t)$ and $L_{1}(T;\alpha^{*})$, as defined in (14), (15), and (19), respectively, in the example systems of the different regimes A,B,C,D. Estimated from numerical simulations.
Building on the above error analysis, we define four parameter regimes of the CNVM, ordered by ascending mMFE error:
  1. 1.
    Regime A: Very similar rates $\big{|}\frac{r_{a,b}}{r_{c,d}}-1\big{|}<0.05$ for all valid combinations $a,b,c,d\in[M]$, and moderate (i.e., small but non-negligible) noise $\tilde{r}_{a,b}$. Example: $r_{0,1}=1$, $r_{1,0}=0.98$, $\tilde{r}_{0,1}=\tilde{r}_{1,0}=0.01$.
  2. 2.
    Regime B: Similar rates $\big{|}\frac{r_{a,b}}{r_{c,d}}-1\big{|}<0.5$ and moderate noise $\tilde{r}_{a,b}$. Example: $r_{0,1}=1$, $r_{1,0}=0.7$, $\tilde{r}_{0,1}=\tilde{r}_{1,0}=0.01$.
  3. 3.
    Regime C: Similar rates and (partially) no noise. Example: $r_{0,1}=1$, $r_{1,0}=0.98$, $\tilde{r}_{0,1}=\tilde{r}_{1,0}=0$.
  4. 4.
    Regime D: Dissimilar rates and (partially) no noise. Example: $r_{0,1}=1$, $r_{1,0}=0$, $\tilde{r}_{0,1}=0$, $\tilde{r}_{1,0}=0.3$ (i.e., SIS model [22]).
(The case of large noise is not considered, as it would render the impact of the underlying network negligible.)
We outline the differences between these regimes using the above-mentioned example parameters (rescaled to exhibit comparable behavior over the interval $t\in[0,100]$). In all regimes, we use the same initial state, in which a randomly chosen $10\%$ of nodes start with opinion 1 and the remainder with opinion 0, and the same BA network with parameter $m=3$ and $N=10^{4}$ nodes. Fig. 1 shows, for each regime, the errors $\delta_{\text{cor}}$ and $\delta_{\text{int}}$ for the optimal $\alpha^{*}$ as defined in (20) with $T=100$, estimated from 15000 model simulations, as well as the resulting $L_{1}$-qerror (19). Fig. 2 (left) shows the corresponding trajectories $\bar{c}(t)$ and $\bar{c}_{\text{{\tiny mMFE}}}(t)$.
Figure 2: Trajectories of $\bar{c}$, estimated from simulations, for the example systems of the different regimes A B,C,D (left), compared to the mMFE (II). The mMFE uses the optimal $\alpha^{*}$ as in (20), shown on the right (green dashed). The right plot also shows the $\alpha$ that minimizes $\delta_{\text{cor}}(t;\alpha)$ for each $t$ (orange solid).
Regime A exhibits the smallest $L_{1}$-error because $|r_{1,0}-r_{0,1}|$ is small and the error terms $\delta_{\text{cor}}$ and $\delta_{\text{int}}$ are close to 0 for all $t$. In regime B the error terms $\delta_{\text{cor}}$ and $\delta_{\text{int}}$ become slightly larger, and since $|r_{1,0}-r_{0,1}|$ also increases compared to regime A, the $L_{1}$-error is slightly larger.
A substantially different behavior appears in regime C. Due to the absence of noise, clusters of identical opinion are highly stable, and the system eventually falls into an absorbing state with $x_{i}=0$ for all $i$ or $x_{i}=1$ for all $i$. As a result, the expected correlation between neighbors increases over time, so that the chosen $\alpha^{*}$ is too small (i.e., $\delta_{\text{cor}}<0$) for $t<20$ but too large (i.e., $\delta_{\text{cor}}>0$) for $t>20$. Since $|\delta_{\text{cor}}|$ is rather large most of the time, the overall $L_{1}$-error is larger than in regimes A and B.
Finally, in regime D (SIS model) the largest error occurs. During the transient phase, up to approximately $t=30$, the infection ($x_{i}=1$) spreads from the initially infected nodes, forming clusters of infected nodes that create strong correlations between neighbors. Over time, these clusters get broken up as nodes spontaneously become susceptible again, leading to an equilibration of $\bar{c}$ and a reduction in neighbor correlations. As a result, $\delta_{\text{cor}}$ is much larger than 0 in the transient phase and well below 0 in equilibrium. Combined with a comparatively large $\delta_{\text{int}}$ and $|r_{1,0}-r_{0,1}|$, the overall $L_{1}$-error is greater than in all other regimes.
The most important observation from this experiment is that in regimes A and B, when imitation rates are similar and noise is present, the mMFE (3) produces only a very small error. In regimes C and D, on the other hand, the correlation between neighbors varies over time, leading to a larger error. The correlation assumption (10) is not sufficiently satisfied in these regimes for a time-constant $\alpha$, as Fig. 2 (right) illustrates.
Figure 3: Interchangeability error $\delta_{\text{int}}$ when using the regular shares $c$ instead of the degree-weighted shares $\bar{c}$ in (15), for regimes A and B (with different initial conditions than in Fig. 1).
It should also be noted that the use of the degree-weighted shares $\bar{c}$, as opposed to the regular shares $c$, is crucial for obtaining a small error in the initial phase. Fig. 3 illustrates that using the regular shares $c$ in the interchangeability assumption (12) instead of $\bar{c}$ can result in an error $\delta_{\text{int}}$ that is magnitudes larger. This is demonstrated using an initial state $x$ where the top $10\%$ of highest-degree nodes have opinion $1$ and the rest opinion 0, resulting in a large discrepancy between $c$ and $\bar{c}$.
II.4 Practical considerations
Computing the factor $\alpha^{*}$ by solving (20) can be computationally demanding as many simulations may be required to estimate the weighted shares trajectory $\bar{c}(t)$, $t\in[0,T]$. However, $\alpha^{*}$ can be calculated efficiently using only the steady-state weighted shares $\bar{c}_{\infty}$, which can be obtained from a single (sufficiently long) simulation started from an arbitrary initial state. Consider that in the steady-state of the mMFE (3) it is
$0=\frac{d}{dt}\bar{c}_{a}=$ (21)
$\alpha\sum_{b\neq a}\bar{c}_{a}\bar{c}_{b}(r_{b,a}-r_{a,b})+\sum_{b\neq a}\bar{c}_{b}\tilde{r}_{b,a}-\bar{c}_{a}\tilde{r}_{a,b}$
for all $a\in[M]$, where summation is performed over all $b$ except $b=a$. This implies that
$\alpha=\frac{\sum_{b\neq a}(\bar{c}_{a}\tilde{r}_{a,b}-\bar{c}_{b}\tilde{r}_{b,a})}{\sum_{b\neq a}\bar{c}_{a}\bar{c}_{b}(r_{b,a}-r_{a,b})}.$ (22)
Hence, $\alpha$ is given by inserting $\bar{c}_{\infty}$ into (22). In the ergodic setting, e.g., if $\tilde{r}_{a,b}>0$ for all $a\neq b$, the above equation yields the optimal $\alpha^{*}$ as defined in (20) with $T\to\infty$.
The usefulness of the mMFE would be severely limited if $\alpha$ had to be calculated from (20) or (22) for every system of interest. However, the next section shows that this is not necessary as the optimal $\alpha$ is robust across initial states and a wide range of system parameters.
III Robustness
The previous section illustrated that, especially in regimes A and B, the mMFE with an appropriate choice of the correlation factor $\alpha$ yields a very good approximation. This section shows that this choice of $\alpha$ is robust to variations in the initial condition and the parameters $r$ and $\tilde{r}$. Thus, the mMFE is general in the sense that a suitable $\alpha$ provides a good approximation for any initial condition and any model parameters in regimes A and B. Moreover, the optimal $\alpha$ depends only on the preferential attachment parameter $m$ of the BA graph, as will be discussed later.
The robustness of $\alpha$ is verified with the following numerical study. We consider a BA network with preferential attachment parameter $m=3$ and $N=10^{5}$ nodes, and the CNVM with $M=3$ opinions. We sample 200 initial states $x\in[M]^{N}$ using four different strategies to obtain a diverse set of random initial conditions:
  1. 1.
    Uniform: The initial opinion of each node is drawn uniformly from $[M]$.
  2. 2.
    Uniform shares: Sample shares $c$ uniformly at random from the simplex $\{c\in[0,1]^{M}\mid\lVert c\rVert_{1}=1\}$, and assign opinions randomly so that $c$ is matched.
  3. 3.
    By degree: Again, sample shares $c$ uniformly. Choose a random order of opinions, assign the first opinion to the $Nc_{1}$ nodes with largest degree, the second opinion to the next $Nc_{2}$ remaining nodes with largest degree, and so on.
  4. 4.
    Clusters: Randomly assign opinions to a few nodes. Then, iteratively, take a random node that already has an opinion and let it pass this opinion to all of its neighbors that do not yet have one. Repeat this process until all nodes have an opinion.
Moreover, for each of the 200 initial states, we sample random parameters from regimes A and B via
$r_{a,b}\sim\text{Uniform}([0.95,1.05])$ (23)
and set $\tilde{r}_{a,b}=0.01$ for all $a,b\in[M]$. The correlation factor $\alpha$ is computed with (22) from a single long simulation starting from a random one of the initial states, which yields
$\alpha=0.81.$ (24)
(In fact, every initial state would result in approximately the same $\alpha$, as this experiment will show.) For each of the 200 initial states and parameter sets, we computed the trajectories of the weighted shares $\bar{c}(t)$ via simulation and compared them to the mMFE solution $\bar{c}_{\text{{\tiny mMFE}}}(t)$ under identical initial conditions and parameters.
The results of this experiment are shown in Fig. 4. We observe errors below 0.005 for all trajectories, with a median error smaller than 0.003. This confirms that our choice of $\alpha=0.81$ yields a very small error for any choice of initial condition and parameter set in regimes A and B, thereby demonstrating the robustness of the mMFE.
Figure 4: $L_{1}$-error (19) of the mMFE for different initial conditions and parameters in regime A and B (top). Even for the two trajectories (a) and (b) with largest errors (bottom) the approximation is very good. Solid lines indicate true weighted shares and dashed/dotted lines the solution of the mMFE.
As mentioned before, the proper choice of $\alpha$ depends only on the preferential attachment parameter $m$. Since $m$ determines the minimum node degree, a larger $m$ generally reduces neighbor correlations. Consequently, as $m$ increases, the optimal $\alpha$ also increases. For very large $m$, neighbors are almost independent, implying $\alpha\approx 1$. This behavior is illustrated in Fig. 5, which shows the optimal $\alpha$ computed from (22) for different values of $m$.
It should be noted that the above results—especially regarding the validity of the mMFE and the robustness of $\alpha$—are accurate only for sufficiently large networks. In the examples discussed here, we found that $N\geq 10^{4}$ is a reasonable threshold. For smaller $N$, the mMFE, like other mean-field approximations, loses accuracy.
Figure 5: The factor $\alpha$ in the mMFE (3) depends on the preferential attachment parameter $m$ of the BA graph.
IV Comparison to classical approximations
Figure 6: The shares $c$ and weighted shares $\bar{c}$ estimated from simulations, in comparison to the mMFE, MFE, and PA. Model parameters from regime A. The left plot shows the trajectories, while the right plot shows the $L_{1}$-error (19) between the mMFE and $\bar{c}$, as well as between MFE and $c$, and PA and $c$.
In this section we compare the mMFE, which is a (M1)(M-1)-dimensional approximation, to other standard low-dimensional moment-closure approximations: the MFE, which also has dimension (M1)(M-1), and the pair-approximation (PA), which has a dimension of (M1)+M(M1)(M-1)+M(M-1). The PA results from closing the moment equations at second order and therefore includes the number of edges between different opinions as additional variables compared to the MFE. For a derivation of the PA, see for example [23]. As Fig. 6 illustrates, the classical approximations (MFE and PA) fail to capture the system behavior on scale-free networks, whereas the mMFE incurs only a very small error. This is not surprising, since the classical approximations assume a homogeneous network structure. The well-known heterogeneous variants of the MFE and PA, or the approximate master equation, are better suited for scale-free networks [24, 4, 15, 25]. However, as these resolve the opinion shares per degree, they involve substantially more variables than the mMFE. For instance, on a BA network with $N=10^{4}$ nodes and $m=4$, there are on average about 100 different node degrees, implying that the heterogeneous MFE has 100 times more variables than our mMFE. Hence, to compare reduced order models with the same degree of efficiency, we only report comparison between the mMFE and the homogeneous MFE and PA.
V Conclusion
In this paper, we derived the mMFE (3), a modified version of the classical mean-field equation (MFE) that describes the dynamics in terms of weighted shares and includes an average “correlation factor” $\alpha$. We then verified that this equation provides an accurate and efficient approximation of the continuous-time noisy voter model on scale-free networks generated by the Barabási–Albert model. It outperforms the standard MFE and pair-approximation (PA), while remaining low-dimensional. We found that the approximation error depends on the underlying parameter regime; in particular, in regimes C and D—corresponding to (partially) noise-free dynamics and/or very dissimilar imitation rates—a time-constant correlation factor $\alpha$ does not exist, causing larger errors.
The inspiration for the mMFE (3) arose from findings in [19], where a computational approach was proposed to identify forecastable variables of dynamical processes on networks. For such processes, reduced modeling is typically guided by theoretical insights—our approach therefore constitutes an exception. Along these lines, it would be highly desirable to gain a theoretical understanding of the observed superior behavior of the mMFE. Furthermore, this work also highlights the importance of data-driven approaches, since many of the insights presented here were supported by computational analysis. These methods could be utilized in future work to derive low-dimensional approximations for other systems that are not yet well understood analytically.
This work in particular could also be extended by refining the mMFE in parameter regimes where its accuracy is limited, for instance by allowing a time-dependent $\alpha$. Moreover, a modified version of the PA that operates on the degree-weighted shares $\bar{c}$ could be explored.
Acknowledgements.
This work has been funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project ID: 546032594.
Appendix A Derivation of mMFE for more than two opinions
In the case of $M>2$ opinions, the degree-weighted shares $\bar{c}\in[0,1]^{M}$ are defined via $\bar{c}=(\bar{c}_{a})_{a\in[M]}$,
$\bar{c}_{a}:=\frac{1}{\sum_{i}d_{i}}\sum_{i\in[N]}d_{i}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}],$ (25)
where $\mathbf{1}_{\{x_{i}=a\}}\in\{0,1\}$ denotes the indicator function of the event $x_{i}=a$. The rate at which node $i$ in state $x_{i}=a$ transitions to a different state $b$ is given by
$r_{a,b}\frac{1}{d_{i}}\sum_{j\in[N]}A_{i,j}\mathbf{1}_{\{x_{j}=b\}}+\tilde{r}_{a,b}=:q_{a,b}^{i}.$ (26)
Hence, for all $a\in[M]$ it is
$\frac{\text{d}}{\text{d}t}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}]$ $=\sum_{b\neq a}\mathbb{P}(x_{i}=b)\ \mathbb{E}[q_{b,a}^{i}\mid x_{i}=b]$ (27)
$\ -\sum_{b\neq a}\mathbb{P}(x_{i}=a)\ \mathbb{E}[q_{a,b}^{i}\mid x_{i}=a].$ (28)
Using the equality $\mathbb{P}(x_{i}=b)\ \mathbb{E}[\mathbf{1}_{\{x_{j}=a\}}\mid x_{i}=b]=\mathbb{E}[\mathbf{1}_{\{x_{j}=a\}}\mathbf{1}_{\{x_{i}=b\}}]$, it follows that
$\frac{\text{d}}{\text{d}t}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}]$ (29)
$=\sum_{b\neq a}\ \frac{r_{b,a}}{d_{i}}\sum_{j}A_{i,j}\mathbb{E}[\mathbf{1}_{\{x_{i}=b\}}\mathbf{1}_{\{x_{j}=a\}}]+\tilde{r}_{b,a}\mathbb{E}[\mathbf{1}_{\{x_{i}=b\}}]$
$\ -\sum_{b\neq a}\ \frac{r_{a,b}}{d_{i}}\sum_{j}A_{i,j}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}\mathbf{1}_{\{x_{j}=b\}}]-\tilde{r}_{a,b}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}].$
Thus, we have
$\frac{\text{d}}{\text{d}t}\bar{c}_{a}=\frac{1}{\sum_{i}d_{i}}\sum_{i}d_{i}\frac{\text{d}}{\text{d}t}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}]=\frac{1}{\sum_{i}d_{i}}\sum_{b\neq a}\Big{(}$ (30)
$r_{b,a}\sum_{i,j}A_{i,j}\mathbb{E}[\mathbf{1}_{\{x_{i}=b\}}\mathbf{1}_{\{x_{j}=a\}}]+\tilde{r}_{b,a}\sum_{i}d_{i}\mathbb{E}[\mathbf{1}_{\{x_{i}=b\}}]$-
$r_{a,b}\sum_{i,j}A_{i,j}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}\mathbf{1}_{\{x_{j}=b\}}]-\tilde{r}_{a,b}\sum_{i}d_{i}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}]\Big{)}.$
Next, we apply the (weak) correlation assumption
$\sum_{i,j}A_{i,j}\mathbb{E}[\mathbf{1}_{\{x_{i}=b\}}\mathbf{1}_{\{x_{j}=a\}}]\overset{!}{=}\\ \sum_{i,j}A_{i,j}\,\alpha\,\mathbb{E}[\mathbf{1}_{\{x_{i}=b\}}]\mathbb{E}[\mathbf{1}_{\{x_{j}=a\}}]$ (31)
and the (weak) interchangeability assumption
$\sum_{i,j}A_{i,j}\mathbb{E}[\mathbf{1}_{\{x_{i}=b\}}]\mathbb{E}[\mathbf{1}_{\{x_{j}=a\}}]\overset{!}{=}\sum_{i,j}A_{i,j}\,\bar{c}_{b}\,\bar{c}_{a}$ (32)
for all $a\neq b$. This yields
$\frac{\text{d}}{\text{d}t}\bar{c}_{a}=\sum_{b\neq a}\Big{(}$ $\alpha\ r_{b,a}\ \bar{c}_{b}\,\bar{c}_{a}+\tilde{r}_{b,a}\,\bar{c}_{b}$ (33)
$-\alpha\ r_{a,b}\ \bar{c}_{a}\,\bar{c}_{b}-\tilde{r}_{a,b}\,\bar{c}_{a}\Big{)}$ (34)
for all $a\in[M]$, which is equivalent to the mMFE (3) when written as a system of equations.